A plane flew 720 miles in 3 hours with a tailwind. It would take 4 hours to travel the same distance with a headwind. Find the speed of the plane in still air and speed of wind.
Answer by Alan3354(69443)

720/3 = 240 mph
720/4 = 180 mph
------------
The airspeed is the average (240 + 180)/2 = 210 mph
The windspeed is the difference = 30 mph